The red color in some fireworks is due to the pres
Answer #1

Sr + 3/2O2 +C(graphite)\rightarrowSrCO3  

1 mole of stontium carbonate formation gives = 1.22*103Kj/mole

147.6 gm of SrCO3 formation released = 1.22*103ki/mole

1.95g of SrCO3 formation released = 1.22*103*1.95/147.6 = 16.11KJ/mole